# Torsion of Rational Elliptic Curves over the Cyclotomic Extensions of $\mathbb{Q}$

### v5

@@ Abstract.
Let $E$ be an elliptic curve defined over $\mathbb{Q}$. In this article, we classify all groups that can arise as $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ up to isomorphism for any prime $p$. When $p-1$ is not divisible by small integers such as $3,4,5,7$, or $11$, we obtain a sharper classification. For any abelian number field $K$, the torsion subgroup $E(K)_{\text{tors}}$ is a subgroup of $E(\mathbb{Q}^{\text{ab}})_{\text{tors}}$. Our methods provide tools to eliminate non-realized torsion structures from the list of possibilities for $E(K)_{\text{tors}}$.
@@ 1. Introduction and Notation
In our work, we classify $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ for any prime $p$. By doing so, we also classify $E(\mathbb{Q}(\mu_{p^{\infty}}))_{\text{tors}}$ for those primes. One of our main results is the following theorem. To derive the shortest possible list, we impose the condition that $p-1$ is not divisible by $3$, $4$, or $5$. The proof is presented in Section 3, while Section 5 examines cases with relaxed conditions to include all primes.
@@
Let $E/\mathbb{Q}$ be an elliptic curve, and let $p>3$ be a prime such that $p-1$ is not divisible by $3$, $4$, or $5$. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s theorem, or one of the following groups:
@@
In the result of González-Jiménez and Najman [10], it is shown that if $K$ is a number field whose degree is not divisible by $2,3,5,$ or $7$, then for any elliptic curve $E/\mathbb{Q}$ we have
@@
We obtain a similar result under the additional assumption that $K/\mathbb{Q}$ is a Galois extension, relaxing the divisibility condition on $2$ to $4$. Similar to their result, we show that many torsion points are either rational points on $E$ or correspond to rational points on a quadratic twist of $E$. This allows us to further restrict the possible torsion group structures using Mazur’s theorem.
@@
This result enables us to classify $E(\mathbb{Q}(\mu_{p^{\infty}}))_{\text{tors}}$ for primes $p$ such that $p-1$ is not divisible by $3,4,5,7,$ or $11$. Furthermore, Gužvić and Vukorepa [13] classified all possible torsion subgroups appearing as $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ for $p=5,7,11$. Our work extends the classification of $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ to all primes $p$. It is important to note that while our classification eliminates most non-realizable torsion subgroups for all primes, some groups in our list may still not be realized for any specific prime $p$.
@@
In several parts of our analysis, we rely on computations performed with SageMath [24]. These codes are publicly available on GitHub at
@@
Acknowledgements. I would like to thank Ekin Ozman for her guidance during my master’s thesis, from which this paper has grown. I am also grateful to Álvaro Lozano-Robledo for his insights on the connection between ramified primes and the field of definition of torsion points. Special thanks to Filip Najman for his helpful discussions, comments, and for bringing related work by his students to my attention. I would also like to express my appreciation to the anonymous referee for their valuable suggestions, particularly the connection between the conductor and quadratic twists, and the insightful idea of using division polynomials to handle all twists simultaneously. Finally, I am deeply thankful to Sueda Senturk Avci for her unwavering support and assistance with computations, especially those involving SageMath.
@@ 2. Torsion of Rational Elliptic Curves
Each of these groups, except for $\mathbb{Z}/15\mathbb{Z}$, appears as the torsion structure over a quadratic field for infinitely many rational elliptic curves $E$. The elliptic curves 50.b3 and 50.a2 have $15$-torsion over $\mathbb{Q}(\sqrt{5})$, 50.b4 and 450.g4 have $15$-torsion over $\mathbb{Q}(\sqrt{-15})$. These are the only rational curves having non-trivial $15$-torsion over any quadratic field.
@@
Let $p\geq 11$, be a prime other than $13$ or $37$. Let $E/\mathbb{Q}$ be an elliptic curve and $P$ be a point of order $p$ on $E$. Then, $[\mathbb{Q}(P):\mathbb{Q}]\geq(p-1)/2$ holds, where $\mathbb{Q}(P)=\mathbb{Q}(x(P),y(P))$ is the extension of rational numbers by the coordinates of the point $P$ called the field of definition of $P$.
@@ 3. Eliminating Possible Torsion
$P^{\sigma}=P$
for all $\sigma\in\operatorname{Gal}(K/\mathbb{Q})$, so $P$ is defined over $\mathbb{Q}$. Since $P$ generates $E(K)[q]$, it follows that
@@
$P^{\sigma^{2}}=P$
for all $\sigma\in\operatorname{Gal}(K/\mathbb{Q})$, which implies
@@
    if the extension degree $n$ is not divisible by $3$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/N\mathbb{Z}$ for $N=14,18,19,21,27,163$, and $E(K)_{\text{tors}}\not\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}$ for $N=7,9$,
@@
    if the extension degree $n$ is not divisible by $3$ and $4$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/N\mathbb{Z}$ for $N=13,37$,
@@
    if the extension degree $n$ is not divisible by $3$ and $7$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/43\mathbb{Z}$,
@@
    if the extension degree $n$ is not divisible by $3$ and $11$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/67\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by $4$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/17\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by $4$, and the fields $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-15})$ are not subfields of $K$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/15\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by $5$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/11\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by $4$ and $5$ then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/25\mathbb{Z}$.
@@
Let us show that $E(K)_{\text{tors}}\cong\mathbb{Z}/N\mathbb{Z}$ is not possible for the values of $N$ other than $14,15,18,21$ given in the lemma, under the appropriate assumptions on the divisors of $n$. All such values of $N$ are either prime or prime powers, so we can apply Lemma 3.1. Under the assumptions on $n$, we have $\gcd(n,\phi(N))\leq 2$, and hence $E(K)_{\text{tors}}\cong E_{d}(\mathbb{Q})_{\text{tors}}$ for some square-free integer $d$ with $\sqrt{d}\in K$. Since $E_{d}$ is also an elliptic curve over $\mathbb{Q}$, Mazur’s theorem implies that $E_{d}(\mathbb{Q})_{\text{tors}}\cong\mathbb{Z}/N\mathbb{Z}$ is not possible.
Showing that $E(K)_{\text{tors}}\cong\mathbb{Z}/N\mathbb{Z}$ is not possible for $N=14$ and $N=18$ can be done similarly, but it requires an additional step. Assume the contrary. Then $E(K)_{\text{tors}}$ has a single $2$-torsion point, which implies that there is a rational $2$-torsion point in $E(K)_{\text{tors}}$. Consequently, for any square-free integer $d$, the torsion subgroup $E_{d}(K)_{\text{tors}}$ also has a rational $2$-torsion point.
If $n$ is not divisible by $3$, then by Lemma 3.1, the torsion subgroup $E_{d}(\mathbb{Q})_{\text{tors}}$ has a point of order $7$ or $9$, respectively, for some square-free integer $d$ with $\sqrt{d}\in K$. Since $E_{d}(\mathbb{Q})_{\text{tors}}$ contains both a $2$-torsion point and a point of order $7$ or $9$, it must contain a point of order $14$ or $18$, respectively. This contradicts Mazur’s theorem.
@@
Similarly, the case $E(K)_{\text{tors}}\cong\mathbb{Z}/15\mathbb{Z}$ also requires special attention. By applying the same arguments, we find that $E_{d_{1}}(\mathbb{Q}(\sqrt{d_{1}d_{2}}))_{\text{tors}}$ has a subgroup of order $15$ for some square-free integers $d_{1},d_{2}$ such that $\sqrt{d_{1}},\sqrt{d_{2}}\in K$. Since $E_{d_{1}}$ is defined over $\mathbb{Q}$, Theorem 2.2 implies that $d_{1}d_{2}=5$ or $d_{1}d_{2}=-15$. However, this is impossible because we assumed both $\sqrt{5}$ and $\sqrt{-15}$ do not belong to the field $K$.
@@
for some square-free integer $d$ with $\sqrt{d}\in K$. Since $E(K)$ has full $2$-torsion, $E_{d}(K)$ also has full $2$-torsion. If the cubic polynomial $x^{3}+Ax+B$ is irreducible over $\mathbb{Q}$, then the degree $n=[K:\mathbb{Q}]$ must be divisible by $3$, contradicting our assumption. Therefore, the cubic must have a rational root, implying that both $E$ and $E_{d}$ have a rational $2$-torsion point. Recall that
@@
Hence, $E_{d}(\mathbb{Q})$ must contain a subgroup of order $2N$, contradicting Mazur’s theorem. ∎
@@
Let $K$ be a finite Galois extension of $\mathbb{Q}$ with $[K:\mathbb{Q}]=n$ for some positive integer $n$. Let $E/\mathbb{Q}$ be an elliptic curve. If the extension degree $n$ is not divisible by $4$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$.
@@
of order $16$. For any $\sigma\in\operatorname{Gal}(K/\mathbb{Q})$, the point $P^{\sigma}\in E(K)_{\text{tors}}$ also has order $16$. Since there are exactly $16$ points of order $16$ in $E(K)_{\text{tors}}$, their $x$-coordinates form a set of $8$ elements, assuming $E$ is given in short Weierstrass form.
@@
This set corresponds to the Galois conjugates of $x_{0}$, and its size equals the degree of the field extension $\mathbb{Q}(x_{0})/\mathbb{Q}$. Since $\mathbb{Q}(x_{0})\subseteq K$ and $[K:\mathbb{Q}]=n$, the degree $[\mathbb{Q}(x_{0}):\mathbb{Q}]$ divides $n$. By assumption, $n$ is not divisible by $4$, so the size of the orbit cannot be $4$ or $8$.
Consider the partitions of $8$ elements into subsets. Either there exists a subset with at most $2$ elements, or the partition consists of two subsets of sizes $3$ and $5$. We will show at the end of the proof that the latter case is not possible.
Without loss of generality, assume that the field extension $\mathbb{Q}(x_{0})$ has degree at most $2$, where $P=(x_{0},y_{0})$ is a point of order $16$. If the orbit size is at most $2$, then $\mathbb{Q}(x_{0})\subseteq\mathbb{Q}(\sqrt{d})$ for some square-free integer $d$. Since
@@
the degree of the field extension $\mathbb{Q}(P)=\mathbb{Q}(x_{0},y_{0})$ divides $4$. Moreover, since the degree also divides $n$, we conclude that $[\mathbb{Q}(P):\mathbb{Q}]\leq 2$.
Now consider the extension $\mathbb{Q}(E[2])$, obtained by adjoining the coordinates of all $2$-torsion points of $E$. Assuming $E$ is given in short Weierstrass form, $\mathbb{Q}(E[2])$ is the splitting field of the cubic polynomial
@@
with $A,B\in\mathbb{Q}$. The point $8P$ is a $2$-torsion point and lies in the field $\mathbb{Q}(P)$. Therefore, one root of the cubic must lie in $\mathbb{Q}(P)$, which is at most quadratic over $\mathbb{Q}$. This implies that the cubic is not irreducible over $\mathbb{Q}$, and hence its splitting field, which is $\mathbb{Q}(E[2])$, has degree at most $2$ over $\mathbb{Q}$.
We have thus shown that there exists a point $P\in E(K)_{\text{tors}}$ of order $16$ such that $\mathbb{Q}(P)$ has degree at most $2$, and that $\mathbb{Q}(E[2])$ also has degree at most $2$. Therefore, the compositum $\mathbb{Q}(E[2],P)$ is an extension of degree dividing $4$, and since it also divides $n$, its degree is at most $2$. This proves that all $2$-torsion points and the point $P$ lie in a quadratic subfield $\mathbb{Q}(\sqrt{d})$ of $K$, for some square-free integer $d$ (possibly $d=1$).
Finally, since all points in $E(K)_{\text{tors}}$ are generated by the $2$-torsion points and a point of order $16$, we conclude that
@@
To conclude our proof, we will show that the set of $x$-coordinates of the $16$-torsion points cannot be partitioned into two Galois orbits of sizes $3$ and $5$ under the action of $\operatorname{Gal}(K/\mathbb{Q})$.
To obtain a contradiction, assume that $P,Q\in E(K)_{\text{tors}}$ are points of order $16$ such that the Galois orbits $\{x_{0}^{\sigma}:\sigma\in\operatorname{Gal}(K/\mathbb{Q})\}$ and $\{x_{1}^{\sigma}:\sigma\in\operatorname{Gal}(K/\mathbb{Q})\}$ contain exactly $3$ and $5$ elements, respectively, where $P=(x_{0},y_{0})$ and $Q=(x_{1},y_{1})$.
This means that $\mathbb{Q}(x_{0})$ is a cubic field and that $\mathbb{Q}(P)$ has extension degree $3$ or $6$. The field $\mathbb{Q}(E[2])$ is the splitting field of the cubic $x^{3}+Ax+B$, so its degree over $\mathbb{Q}$ is $1$, $2$, $3$, or $6$. Moreover, since all of $E(K)_{\text{tors}}$ is generated by $P$ and the $2$-torsion points, we have
@@
Since $n$ is not divisible by $4$, the degree $[\mathbb{Q}(L):\mathbb{Q}]$ is also not divisible by $4$. Moreover, it must divide $36$. Hence, we must have
@@
As a combined result of Lemma 3.1, by assuming that $p-1$ is not divisible by $3,4,5,7,11,$ we can use all items of the lemma except the result about $\mathbb{Z}/15\mathbb{Z}$. This result requires more conditions. We know that when $p$ is a prime, $\mathbb{Q}(\zeta_{p})$ has a unique quadratic subfield. If $p\equiv 1\pmod{4}$, that unique quadratic subfield is $\mathbb{Q}(\sqrt{p})$. If $p\equiv 3\pmod{4}$, as in our theorem, the unique quadratic subfield is $\mathbb{Q}(\sqrt{-p})$. This means that $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-15})$ are not subfields of $\mathbb{Q}(\zeta_{p})$, so we can use that result of the lemma too.
@@
under the additional assumptions that $p-1$ is not divisible by $7$ or $11$. Next, we demonstrate how these additional assumptions can be removed.
If we remove the conditions that $p-1$ is not divisible by $7$ and $11$, then we cannot exclude the possibility of $\mathbb{Z}/43\mathbb{Z}$ or $\mathbb{Z}/67\mathbb{Z}$ using Lemma 3.1 alone. Thus, we need a different strategy. Let
@@
By Lemma 4.7, this is only possible when $p=43$ and $p=67$, respectively, but both cases are excluded by our assumption that $p-1$ is not divisible by $3$.
There is also another way to see that $\mathbb{Z}/43\mathbb{Z}$ or $\mathbb{Z}/67\mathbb{Z}$ cannot be realized over $\mathbb{Q}(\zeta_{p})$ under the assumption that $p-1$ is not divisible by $3$. We use a similar idea to the proof of Lemma 3.2. Let $P\in E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ be a point of order $\ell$ for $\ell=43$ or $\ell=67$. For any $\sigma\in\operatorname{Gal}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$, consider the Galois action. Then $\sigma^{(\ell-1)/3}$ fixes any $\ell$-order point in $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ because

### v6

@@ Abstract.
Let $E$ be an elliptic curve defined over $\mathbb{Q}$. In this article, we classify all groups that can arise as $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ up to isomorphism for any prime $p$. When $p-1$ is not divisible by small integers such as 3,4,5,73,4,5,73 , 4 , 5 , 7, or 111111, we obtain a sharper classification. For any abelian number field $K$, the torsion subgroup $E(K)_{\text{tors}}$ is a subgroup of $E(\mathbb{Q}^{\text{ab}})_{\text{tors}}$. Our methods provide tools to eliminate non-realized torsion structures from the list of possibilities for $E(K)_{\text{tors}}$.
@@ 1. Introduction and Notation
In our work, we classify $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ for any prime $p$. By doing so, we also classify $E(\mathbb{Q}(\mu_{p^{\infty}}))_{\text{tors}}$ for those primes. One of our main results is the following theorem. To derive the shortest possible list, we impose the condition that $p-1$ is not divisible by 333, 444, or 555. The proof is presented in Section 3, while Section 5 examines cases with relaxed conditions to include all primes.
@@
Let $E/\mathbb{Q}$ be an elliptic curve, and let $p>3$ be a prime such that $p-1$ is not divisible by 333, 444, or 555. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s theorem, or one of the following groups:
@@
In the result of González-Jiménez and Najman [10], it is shown that if $K$ is a number field whose degree is not divisible by 2,3,5,2,3,5,2 , 3 , 5 , or 777, then for any elliptic curve $E/\mathbb{Q}$ we have
@@
We obtain a similar result under the additional assumption that $K/\mathbb{Q}$ is a Galois extension, relaxing the divisibility condition on 222 to 444. Similar to their result, we show that many torsion points are either rational points on $E$ or correspond to rational points on a quadratic twist of $E$. This allows us to further restrict the possible torsion group structures using Mazur’s theorem.
@@
This result enables us to classify $E(\mathbb{Q}(\mu_{p^{\infty}}))_{\text{tors}}$ for primes $p$ such that $p-1$ is not divisible by 3,4,5,7,3,4,5,7,3 , 4 , 5 , 7 , or 111111. Furthermore, Gužvić and Vukorepa [13] classified all possible torsion subgroups appearing as $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ for $p=5,7,11$. Our work extends the classification of $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ to all primes $p$. It is important to note that while our classification eliminates most non-realizable torsion subgroups for all primes, some groups in our list may still not be realized for any specific prime $p$.
@@
In several parts of our analysis, we rely on computations performed with SageMath. These codes are publicly available on GitHub at
@@
Acknowledgements. I would like to thank Ekin Ozman for her guidance during my master’s thesis, from which this paper has grown. I am also grateful to Бlvaro Lozano-Robledo for his insights on the connection between ramified primes and the field of definition of torsion points. Special thanks to Filip Najman for his helpful discussions, comments, and for bringing related work by his students to my attention. I would also like to express my appreciation to the anonymous referee for their valuable suggestions, particularly the connection between the conductor and quadratic twists, and the insightful idea of using division polynomials to handle all twists simultaneously. I am additionally thankful to the referee for pointing out the method of factoring over local fields to obtain faster results. Finally, I am deeply thankful to Sueda Senturk Avci for her unwavering support and assistance with computations, especially those involving SageMath.
@@ 2. Torsion of Rational Elliptic Curves
Each of these groups, except for $\mathbb{Z}/15\mathbb{Z}$, appears as the torsion structure over a quadratic field for infinitely many rational elliptic curves $E$. The elliptic curves 50.b3 and 50.a2 have 151515-torsion over $\mathbb{Q}(\sqrt{5})$, 50.b4 and 450.g4 have 151515-torsion over $\mathbb{Q}(\sqrt{-15})$. These are the only rational curves having non-trivial 151515-torsion over any quadratic field.
@@
Let $p\geq 11$, be a prime other than 131313 or 373737. Let $E/\mathbb{Q}$ be an elliptic curve and $P$ be a point of order $p$ on $E$. Then, $[\mathbb{Q}(P):\mathbb{Q}]\geq(p-1)/2$ holds, where $\mathbb{Q}(P)=\mathbb{Q}(x(P),y(P))$ is the extension of rational numbers by the coordinates of the point $P$ called the field of definition of $P$.
@@ 3. Eliminating Possible Torsion
$P^{\sigma}=P\quad\text{for all }\sigma\in\operatorname{Gal}(K/\mathbb{Q}),$
so $P$ is defined over $\mathbb{Q}$. Since $P$ generates $E(K)[q]$, it follows that
@@
$P^{\sigma^{2}}=P\quad\text{for all }\sigma\in\operatorname{Gal}(K/\mathbb{Q}),$
which implies
@@
    if the extension degree $n$ is not divisible by 333, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/N\mathbb{Z}$ for $N=14,18,19,21,27,163$, and $E(K)_{\text{tors}}\not\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}$ for $N=7,9$,
@@
    if the extension degree $n$ is not divisible by 333 and 444, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/N\mathbb{Z}$ for $N=13,37$,
@@
    if the extension degree $n$ is not divisible by 333 and 777, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/43\mathbb{Z}$,
@@
    if the extension degree $n$ is not divisible by 333 and 111111, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/67\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by 444, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/17\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by 444, and the fields $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-15})$ are not subfields of $K$, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/15\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by 555, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/11\mathbb{Z}$.
@@
    if the extension degree $n$ is not divisible by 444 and 555 then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/25\mathbb{Z}$.
@@
Let us show that $E(K)_{\text{tors}}\cong\mathbb{Z}/N\mathbb{Z}$ is not possible for the values of $N$ other than 14,15,18,2114,15,18,2114 , 15 , 18 , 21 given in the lemma, under the appropriate assumptions on the divisors of $n$. All such values of $N$ are either prime or prime powers, so we can apply Lemma 3.1. Under the assumptions on $n$, we have $\gcd(n,\phi(N))\leq 2$, and hence $E(K)_{\text{tors}}\cong E_{d}(\mathbb{Q})_{\text{tors}}$ for some square-free integer $d$ with $\sqrt{d}\in K$. Since $E_{d}$ is also an elliptic curve over $\mathbb{Q}$, Mazur’s theorem implies that $E_{d}(\mathbb{Q})_{\text{tors}}\cong\mathbb{Z}/N\mathbb{Z}$ is not possible.
Showing that $E(K)_{\text{tors}}\cong\mathbb{Z}/N\mathbb{Z}$ is not possible for $N=14$ and $N=18$ can be done similarly, but it requires an additional step. Assume the contrary. Then $E(K)_{\text{tors}}$ has a single 222-torsion point, which implies that there is a rational 222-torsion point in $E(K)_{\text{tors}}$. Consequently, for any square-free integer $d$, the torsion subgroup $E_{d}(K)_{\text{tors}}$ also has a rational 222-torsion point.
If $n$ is not divisible by 333, then by Lemma 3.1, the torsion subgroup $E_{d}(\mathbb{Q})_{\text{tors}}$ has a point of order 777 or 999, respectively, for some square-free integer $d$ with $\sqrt{d}\in K$. Since $E_{d}(\mathbb{Q})_{\text{tors}}$ contains both a 222-torsion point and a point of order 777 or 999, it must contain a point of order 141414 or 181818, respectively. This contradicts Mazur’s theorem.
@@
Similarly, the case $E(K)_{\text{tors}}\cong\mathbb{Z}/15\mathbb{Z}$ also requires special attention. By applying the same arguments, we find that $E_{d_{1}}(\mathbb{Q}(\sqrt{d_{1}d_{2}}))_{\text{tors}}$ has a subgroup of order 151515 for some square-free integers $d_{1},d_{2}$ such that $\sqrt{d_{1}},\sqrt{d_{2}}\in K$. Since $E_{d_{1}}$ is defined over $\mathbb{Q}$, Theorem 2.2 implies that $d_{1}d_{2}=5$ or $d_{1}d_{2}=-15$. However, this is impossible because we assumed both $\sqrt{5}$ and $\sqrt{-15}$ do not belong to the field $K$.
@@
for some square-free integer $d$ with $\sqrt{d}\in K$. Since $E(K)$ has full 222-torsion, $E_{d}(K)$ also has full 222-torsion. If the cubic polynomial $x^{3}+Ax+B$ is irreducible over $\mathbb{Q}$, then the degree $n=[K:\mathbb{Q}]$ must be divisible by 333, contradicting our assumption. Therefore, the cubic must have a rational root, implying that both $E$ and $E_{d}$ have a rational 222-torsion point. Recall that
@@
[…]
@@
Let $K$ be a finite Galois extension of $\mathbb{Q}$ with $[K:\mathbb{Q}]=n$ for some positive integer $n$. Let $E/\mathbb{Q}$ be an elliptic curve. If the extension degree $n$ is not divisible by 444, then $E(K)_{\text{tors}}\not\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$.
@@
of order 161616. For any $\sigma\in\operatorname{Gal}(K/\mathbb{Q})$, the point $P^{\sigma}\in E(K)_{\text{tors}}$ also has order 161616. Since there are exactly 161616 points of order 161616 in $E(K)_{\text{tors}}$, their $x$-coordinates form a set of 888 elements, assuming $E$ is given in short Weierstrass form.
@@
This set corresponds to the Galois conjugates of $x_{0}$, and its size equals the degree of the field extension $\mathbb{Q}(x_{0})/\mathbb{Q}$. Since $\mathbb{Q}(x_{0})\subseteq K$ and $[K:\mathbb{Q}]=n$, the degree $[\mathbb{Q}(x_{0}):\mathbb{Q}]$ divides $n$. By assumption, $n$ is not divisible by 444, so the size of the orbit cannot be 444 or 888.
Consider the partitions of 888 elements into subsets. Either there exists a subset with at most 222 elements, or the partition consists of two subsets of sizes 333 and 555. We will show at the end of the proof that the latter case is not possible.
Without loss of generality, assume that the field extension $\mathbb{Q}(x_{0})$ has degree at most 222, where $P=(x_{0},y_{0})$ is a point of order 161616. If the orbit size is at most 222, then $\mathbb{Q}(x_{0})\subseteq\mathbb{Q}(\sqrt{d})$ for some square-free integer $d$. Since
@@
the degree of the field extension $\mathbb{Q}(P)=\mathbb{Q}(x_{0},y_{0})$ divides 444. Moreover, since the degree also divides $n$, we conclude that $[\mathbb{Q}(P):\mathbb{Q}]\leq 2$.
Now consider the extension $\mathbb{Q}(E[2])$, obtained by adjoining the coordinates of all 222-torsion points of $E$. Assuming $E$ is given in short Weierstrass form, $\mathbb{Q}(E[2])$ is the splitting field of the cubic polynomial
@@
[…]
We have thus shown that there exists a point $P\in E(K)_{\text{tors}}$ of order 161616 such that $\mathbb{Q}(P)$ has degree at most 222, and that $\mathbb{Q}(E[2])$ also has degree at most 222. Therefore, the compositum $\mathbb{Q}(E[2],P)$ is an extension of degree dividing 444, and since it also divides $n$, its degree is at most 222. This proves that all 222-torsion points and the point $P$ lie in a quadratic subfield $\mathbb{Q}(\sqrt{d})$ of $K$, for some square-free integer $d$ (possibly $d=1$).
Finally, since all points in $E(K)_{\text{tors}}$ are generated by the 222-torsion points and a point of order 161616, we conclude that
@@
To conclude our proof, we will show that the set of $x$-coordinates of the 161616-torsion points cannot be partitioned into two Galois orbits of sizes 333 and 555 under the action of $\operatorname{Gal}(K/\mathbb{Q})$.
To obtain a contradiction, assume that $P,Q\in E(K)_{\text{tors}}$ are points of order 161616 such that the Galois orbits $\{x_{0}^{\sigma}:\sigma\in\operatorname{Gal}(K/\mathbb{Q})\}$ and $\{x_{1}^{\sigma}:\sigma\in\operatorname{Gal}(K/\mathbb{Q})\}$ contain exactly 333 and 555 elements, respectively, where $P=(x_{0},y_{0})$ and $Q=(x_{1},y_{1})$.
This means that $\mathbb{Q}(x_{0})$ is a cubic field and that $\mathbb{Q}(P)$ has extension degree 333 or 666. The field $\mathbb{Q}(E[2])$ is the splitting field of the cubic $x^{3}+Ax+B$, so its degree over $\mathbb{Q}$ is 111, 222, 333, or 666. Moreover, since all of $E(K)_{\text{tors}}$ is generated by $P$ and the 222-torsion points, we have
@@
Since $n$ is not divisible by 444, the degree $[\mathbb{Q}(L):\mathbb{Q}]$ is also not divisible by 444. Moreover, it must divide 363636. Hence, we must have
@@
As a combined result of Lemma 3.1, by assuming that $p-1$ is not divisible by 3,4,5,7,11,3,4,5,7,11,3 , 4 , 5 , 7 , 11 , we can use all items of the lemma except the result about $\mathbb{Z}/15\mathbb{Z}$. This result requires more conditions. We know that when $p$ is a prime, $\mathbb{Q}(\zeta_{p})$ has a unique quadratic subfield. If $p\equiv 1\pmod{4}$, that unique quadratic subfield is $\mathbb{Q}(\sqrt{p})$. If $p\equiv 3\pmod{4}$, as in our theorem, the unique quadratic subfield is $\mathbb{Q}(\sqrt{-p})$. This means that $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-15})$ are not subfields of $\mathbb{Q}(\zeta_{p})$, so we can use that result of the lemma too.
@@
under the additional assumptions that $p-1$ is not divisible by 777 or 111111. Next, we demonstrate how these additional assumptions can be removed.
If we remove the conditions that $p-1$ is not divisible by 777 and 111111, then we cannot exclude the possibility of $\mathbb{Z}/43\mathbb{Z}$ or $\mathbb{Z}/67\mathbb{Z}$ using Lemma 3.1 alone. Thus, we need a different strategy. Let
@@
By Lemma 4.7, this is only possible when $p=43$ and $p=67$, respectively, but both cases are excluded by our assumption that $p-1$ is not divisible by 333.
There is also another way to see that $\mathbb{Z}/43\mathbb{Z}$ or $\mathbb{Z}/67\mathbb{Z}$ cannot be realized over $\mathbb{Q}(\zeta_{p})$ under the assumption that $p-1$ is not divisible by 333. We use a similar idea to the proof of Lemma 3.2. Let $P\in E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ be a point of order $\ell$ for $\ell=43$ or $\ell=67$. For any $\sigma\in\operatorname{Gal}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$, consider the Galois action. Then $\sigma^{(\ell-1)/3}$ fixes any $\ell$-order point in $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ because